A car starts from rest in the pit area and accelerates at a uniform rate to a speed of 35 m/s in 11 s, moving on a circular track of radius 500 m. Determine the tangential and centripetal components of the net force exerted on a car by the ground when is speed is 15 m/s. The car’s mass is 950 kg. (Assume constant tangential acceleration).
Answer #1

a)

From Kinematic equation

V=Vo+at

Since it starts from rest Vo=0

Therefore Tangential acceleration

35=0+at*11

at=35/11 =3.182 m/s2

Tangential force

Ft = ma = 950*3.182 =3022.73 N

b)

Centripetal acceleration is

ac=V2/r =152/500 =0.45 m/s2

Centripetal fore is given by

Fc=ma =950*0.45 =427.5 N